OpenStudy (anonymous):

The lengths of the legs of a right triangle are 15 cm and 36 cm. What is the length of the hypotenuse? A. 57 centimeters B. 21 centimeters C. 51 centimeters D. 39 centimeters

OpenStudy (anonymous):

do i use A^+b^=C^?

OpenStudy (jjuden):

yes

OpenStudy (jjuden):

a^2+b^2=C^2

OpenStudy (anonymous):

225+1296=C^2?

OpenStudy (jjuden):

yup

OpenStudy (anonymous):

and i just add it? i got 1521

OpenStudy (jjuden):

now that you have 1521 go though each answer and multiply it by it's self untill you end you with 1521

OpenStudy (anonymous):

its 39 right?

OpenStudy (dtan5457):

Yeah. \[\sqrt{1521}=39\]

OpenStudy (jjuden):

yes correct

OpenStudy (anonymous):

Thanks you guys!

OpenStudy (jjuden):

welcome